OpenStudy (anonymous):

Lance Armstrong wants to travel from Orlando to Tampa (117 miles) on his bike in 180 minutes or less. What speed must he travel to achieve his goal? 65 miles per hour 39 miles per hour 15 miles per hour 26 miles per hour

OpenStudy (anonymous):

39 mph

OpenStudy (anonymous):

65 miles per hour

OpenStudy (anonymous):

2 different answers....

OpenStudy (anonymous):

it is 39 soz i did it wrong the firsdt time xD

OpenStudy (anonymous):

117 miles = rate * 180 min rate = 117/180 miles / min rate = .65 miles / min = 39 miles / hour

OpenStudy (anonymous):

thanks :)